OpenStudy (anonymous):

A school is having a canned food drive. Each class is challenged to collect 140 cans. If there are c classes in the school, which equation could be used to find the total number of cans, T, the school will collect if each class meets the challenge?

OpenStudy (anonymous):

how many classes and students are there and how many do each student bring

OpenStudy (anonymous):

\[T =140c , T= 140divc , T= c -140, T= 140+c\]

OpenStudy (anonymous):

ok so say there is 5 classes with 20 students in each they need to make 140 cans how many will each student bring